OpenStudy (anonymous):

The formula below gives A, the area of a circle with radius r. A(r) = 3.14 r2 To the nearest hundredth, what is the value of A(6)? (Use 3.14 for .) A. 37.68 B. 18.84 C. 113.04 D. 354.95

OpenStudy (zale101):

\[A(r) = 3.14 *r^2 \] then \(A(6) = 3.14 *6^{2}\) =